OpenStudy (arkay):

let f(x)=-2x+4 and g(x)=-6x-7. find f(x)-g(x)

OpenStudy (luffingsails):

Here's an example of this. f(x) = 7 and g(x) = 4, find f(x) - g(x) answer: f(x) - g(x) = 7 - 4 = 3 Just replace what you know f(x) and g(x) to be and then simplify.

OpenStudy (arkay):

i got 4x+11

OpenStudy (triciaal):

f(x) -g(x) (-2x + 4) - (-6 x -7) -2x +6x +4 + 7 correct